find 3 consecutive multiples of 5 whose sum is 60
Answer by 303795(562)

Let the first number be x. The second multiple of 5 will be x+5 and the third will be x + 10.
These add to give x+ x + 5 + x + 10 = 3x + 15
3x + 15 = 60
3x = 45
x = 15.
The second number will be 20 and the third 25 (15 + 20 + 25 = 60)